11:15PM z-4: cat test_error                ~/programs/ocaml/bound [jl/ttypts/4]
test_examples 800
test_errors 192
delta 0.025
lower_delta 0.025
11:15PM z-5: bound test_error              ~/programs/ocaml/bound [jl/ttypts/4]
Applying varying approximation tail bound
delta 0.025
lower_delta 0.025
test_examples 800
test_errors 192
approximation automatic
true_error = 0.24 0.28252880089 0.20065891277